Please help me solve this. Find the equation of the parabola determined by the given information. Focus is (0,- sqrt (7)) and the directrix is y= sqrt (7)
Answer by lwsshak3(11628)

Please help me solve this. Find the equation of the parabola determined by the given information. Focus is (0,- sqrt (7)) and the directrix is y= sqrt (7)
This is a parabola that opens downward.
Its standard form of equation: (x-h)^2=-4p(y-k)
vertex:(0,0)
axis of symmetry: x=0 or y-axis
p=√7 (distance from vertex to directrix and to focus on the axis of symmetry)
4p=4√7
equation of parabola: x^2=-4√7 y